is 4 0% According to the dividend growth model, what is this stocks total expected rate of return? yr Cop paditsannualdidend of $2 30 per share and today you wish to purchase the stock at todays quoted price of $27 49. You beleve that the dividend growth rate 0 115% 10 5% 95% @ 140% 0.127%
Answer #1

As per Dividend growth model formula is,  

Price = D1/Re -g

27.49 = 2.3*1.04/Re - .04

27.49 *Re - 27.49* .04 = 2.392

27.49Re - 1.0996 = 2.392

Re = 2.392 + 1.0996/27.49

= 3.4916/27.49 * 100

12.7 %

Therefore expected Return = 12.7%